Price of a book increases 15% successively (2times) what is the new price of the book more compared to that of the old price

Grade:11

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Student

Suppose the initial price is 100

after first increment price will be 100+15=115

2nd increment will be calculated on 115

15% of 115 will be=.15x115=17.25

new price=115+17.25=132.25

total increment=32.35

Increment %=32.25/100 = 32.25%

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..